Madam Oak’s Furniture manufactures wooden rocking chairs. Madam Oak’s Furniture purchased the following materials in June: 1,500 springs (part of the rocking mechanism) at a cost of $15,000; each chair uses 2 springs Glue at a cost of $1,500 (enough to manufacture 500 chairs) Stain at a cost of $500 (enough to manufacture 500 chairs) Wood at a cost of $5,000 (enough to build 1,000 chair frames) Madam Oak’s Furniture produced 500 chairs during June and had no beginning balances in raw materials, work in process, or finished goods inventories. Required A. Which of Madam Oak’s material costs would likely be classified as direct materials? Which as indirect materials? B. If Madam Oak’s Furniture completed only 400 of the 500 chairs on which production was started, what is the finished goods inventory account balance at the end of June? C. If 380 of the finished chairs were sold, what would be the cost of goods sold? D. What is the work in process inventory account balance at the end of June?
Answer #1

Given,

1500 spring. = $15000. (2 springs for each chair)

Glue. = $1500. (For 500 chairs)

Stain. = $500. (For 500 chairs)

Wood. = $5000. ( for 1000 chairs)

Production is 500 chairs, no opening inventory.

Required A,

Direct material for 500 chairs

Spring = $10000

( 2 spring × 500 chairs= 1000spring).

Glue. = $1500

Stain. = $500

Wood. = $2500

(5000÷1000×500)

::total. $14500

Indirect material

Spring = $ 5000

($15000-$10000)

Wood. = $2500

($5000-$2500)

:: total. $7500

Required B,

Cost of finished goods inventory for 400 chairs,

Spring(15000÷1500×800). = $8000

(400× 2 spring). = 800 springs

Glue (1500÷500×400). = $1200

Stain (500÷500×400). =$ 400

Wood(5000÷1000×400). =$2000

:: total. = $11600

Required C,

Cost of 380 chairs sold:

Spring (15000÷1500×760). = $7600

(380 chairs×2 springs)

Glue(1500÷500×380). = $1140

Stain (500÷500×380). = $380

Wood(5000÷1000×380). = $1900

::total. = $11020

Required D,

Total chairs. = 500

Chairs produced = 400

Therefore,

Work in progress. = 100

Cost of work in process(100 chairs)

Spring (15000÷1500×200). = $2000

(100 chairs × 2 spring)

Glue (1500÷500×100). = $300

Stain(500÷500×100). = $100

Wood (5000÷1000×100). =$500

::total. = $2900

Summary,

Determination of cost of goods sold,work in progress.